# On the diversity of strongly-interacting Type IIn supernovae

### v1

@@ Abstract
Context. Massive stars experience strong mass-loss, producing a dense, H-rich circumstellar medium (CSM). After the explosion, the collision and continued interaction of the supernova (SN) ejecta with the CSM power the light curve through the conversion of kinetic energy into radiation. When the interaction is strong, the light curve shows a broad peak and high luminosity lasting for a relatively long time. Also the spectral evolution is slower, compared to non-interacting SNe. Energetic shocks between the ejecta and the CSM create the ideal conditions for particle acceleration and production of high-energy (HE) neutrinos above 1 TeV.
@@
The net result of an enormous (, Kiewe et al. 2012) mass-loss is a massive circumstellar medium (CSM) that is revealed by narrow lines in the spectrum after the explosion (e.g., Type IIn SNe, Schlegel 1990, 1996; Filippenko 1997) and contributes to the luminosity through interaction with the SN ejecta. The luminosity of interacting SNe is mainly powered by the conversion of kinetic energy into radiation in the shock of the ejecta with the CSM. It is expected that the shock is stronger when a shell ejection occurs shortly before the explosion (less than a few years). In fact, the closer the CSM shell is to the progenitor, the denser it is, and thus the SN ejecta colliding into it give rise to a more intense shock wave. When the CSM has a high density, the interaction can completely mask the internal power source to the extent that the underlying event can even be a non-terminal outburst (Vink, 2015) rather than a CC SN, or even a thermonuclear explosion (Silverman et al., 2013). The density of the CSM and its spatial distribution affect the shape of the light curve (Khatami & Kasen, 2023) and the strength of the emission lines, giving rise to asymmetries (Andrews et al., 2017) and peculiar line profiles such as double peaks (Andrews & Smith, 2018). Not only that, the mass and density of the CSM can also affect the shape of the light curve (Khatami & Kasen, 2023). Interestingly, Taddia et al. (2013) find that there does not appear to be a continuity of properties among SNe IIn but rather different subtypes, possibly hinting at the presence of different progenitors.
A well-known SN that showed huge interaction is SN 2010jl (Smith et al., 2012b; Fransson et al., 2014), a luminous Type IIn SN with a bright ( mag at peak) light curve that lasted more than 1000 and spectra dominated by H Balmer lines with symmetric, electron-scattering driven profiles. These characteristics are interpreted as signs of interaction with a massive (), H-rich CSM, possibly produced by an LBV progenitor (Fransson et al., 2014). Another example is SN 2013L (Andrews et al., 2017; Taddia et al., 2020), which also showed a prolonged light curve lasting 1500 and peaking at around mag. The spectra are also slow-evolving and dominated by Balmer lines but the profile of the H line is significantly asymmetric, with a blue shoulder that is interpreted as due to the emergence of the shock wave not fully hidden by electron scattering (Taddia et al., 2020). We will use these two SNe as reference throughout the paper.
@@
Interaction is a powerful phenomenon that was proposed to power many SNe, among which, the superluminous SN class (SLSNe, Smith et al. 2007; Gal-Yam 2012), so as to explain why they are so luminous and with such slow evolution. The mechanism powering SLSNe is still debated, but the slow rise seems to favour a central engine (particularly, a magnetar, Kasen & Bildsten 2010) in a massive ejecta with long diffusion time.
In some cases, the CSM distribution may be strongly asymmetric. If it is also particularly close to the progenitor, the ejecta can quickly and completely engulf it. Narrow lines are then no longer visible in the spectra and the broad lines only show P-Cygni profiles typical of an expanding photosphere, but interaction continues within the inner ejecta, providing energy to the light curve. In these cases, it is difficult to determine whether the luminosity is fully due to interaction or to a central engine such as a magnetar, but the extreme energy and duration of these events seems to point toward a combination of the two phenomena (Kangas et al., 2022; Pessi et al., 2023; Salmaso et al., 2023).
@@ 2 The sample
The sample was built selecting newly-discovered transients announced through services such as AstroNotes111https://www.wis-tns.org/astronotes and Astronomer’s Telegrams222https://astronomerstelegram.org that met our selection criteria. We evaluated both the classification spectrum and the light curve, and selected SNe classified as Type IIn, with relatively fast rising time ( days), and high but not extreme luminosity (). These criteria tend to exclude SLSNe, which are at least one order of magnitude more luminous ( or mag) and tend to have longer rise time to peak ( d) (Lunnan et al., 2015). Also, a preference was given to nearby events (typically within redshift , so that a more complete follow-up could be ensured). Following our procedure, we selected 11 targets in three years, which are listed in Tab. 3, together with their coordinates, redshift, distance modulus, and galactic absorption. We also report some notable parameters derived from the pseudo-bolometric light curve (for details on the computation, see Sect. 3), namely: the rise time , defined as the time between the last non-detection and the first peak; the peak luminosity ; the integrated luminosity in the first 200 days . The case of SN 2021adxl is a bit peculiar since its poorly-constrained rise time is days, but given the light curve shape and luminosity, very similar to SN 2010jl (Fransson et al., 2014), it is likely that the explosion date was shortly before the peak. Therefore, we elected to keep it in the sample.
@@
As we will detail in the following, it turned out that not all SNe we selected with our criteria proved to evolve as expected and 7 out of 11 candidates were excluded from more detailed analysis for not showing signs of strong interaction. Their light curve was generally dimmer than the four strongly-interacting SNe that remained in the sample, and faded faster. This can clearly be seen in Fig. 1, where we show their pseudo-bolometric light curve built using only optical filters (see Sect. 3 for the computation).
@@
To discriminate the transients that display strong interaction, we used the measured luminosity integrated over the first 200 and imposed that this parameter . This criterion helps identify the objects with prolonged, strong interaction. In fact, with longer integration times, the contamination arises from objects undergoing prolonged but mild interaction (). Conversely, with shorter integration times there is significant contamination from objects experiencing strong interaction for a limited period (e.g., SN 1998S). Our parameter is a compromise between the two cases and ensures that all the selected SNe are strongly-interacting. The excluded SNe had a lower energy display in the first 200 days compared to the selected ones. This can be seen from Fig. 2, where we also add the values calculated for the strongly-interacting SN 2010jl, the mildly-interacting SN 1998S, and the normal Type IIP SN 2004et Sahu et al. (2006) as reference for different amounts of interaction. An intermediate case is that of SN 2021gci, which has significantly more energy than the others, and with a more luminous first peak, albeit still one order of magnitude below the selected transients. On the other hand, SNe 2022iaz and 2022owx have more luminous peaks but faded too fast, providing a small total amount of energy.
@@
This exercise denotes that a single spectrum and the early light curve is not enough to discern between strongly-interacting and normal interacting SNe, thus causing the sample to be contaminated with objects that have to be rejected later. The very minimum information required to identify a strongly-interacting SN include at least photometric coverage of 2/3 months and a couple of good resolution spectra with enough signal-to-noise (S/N) at different epochs to check the evolution. Ideally, one should have a well-sampled light curve for the first 200 days to properly distinguish between strongly-interacting SNe and SNe with low interaction based on the energy input.
@@
The four SNe selected here that fulfill our criteria are then 2021acya, 2021adxl, 2022qml, and 2022wed. In the following, we will describe each object in more detail. We note that a preprint presenting detailed observations for one of the targets in our list, SN 2021adxl, was posted in the open access archive arXiv (Brennan et al., 2023, now in press). We will refer to their observations and results in the relevant sections.
@@ 2.1 SN 2021acya
SN 2021acya was discovered on October 2021 (MJD 59518.029) in the band by ATLAS at a magnitude of 18.124, while the last non-detection was only two days prior (Tonry et al., 2021). As the explosion epoch, then, we took MJD . It was then classified as SN IIn on November 2021 (Ragosta et al., 2021).
@@ 2.3 SN 2022qml
The spectral evolution of SN 2022qml is shown in Fig. 7. The spectrum at +24 has a blue continuum, implying a high temperature, and the only notable feature is the narrow H on top of a very broad but shallow emission. The spectrum at +49 starts to show a distinctive blue bump due to a forest of Fe II lines and a broad, asymmetric H with a boxy profile smoothed by electron scattering wings. We also identify the narrow line of [Fe X] and, possibly, [Ne V] , [Fe XI] and a broader emission that could be due to a blend of coronal [Fe IV] with [Ca V] . The absorption in the P-Cygni profile of the narrow H is also clearly visible, which allows for measuring the velocity of the progenitor wind (). This feature is not identifiable in the other spectra of this object due to their lower resolution. At +56 the only prominent features are Balmer lines, [Fe X] , and He I . After that, the high ionisation lines disappear while the spectra remain dominated by the H, until phase +111, when the Ca II NIR emerges with a broad, symmetric emission. A zoom on the H region of this spectrum is also plotted in Fig. 6. The shape of the blue shoulder is similar to SNe 2021adxl and 2013L but it is broader, indicating a higher shock velocity.
The strong blue bump of SN 2022qml is reminiscent of a feature seen in Type Ia-CSM SNe (e.g., SN 2002ic, Kotak et al. 2004) but also in Type Ic SNe (e.g., SN 1998bw Galama et al. 1998; Kulkarni et al. 1998). As we mentioned, the presence of the blue bump is usually attributed to a forest of Fe II lines. In the case of SN 2022qml, the strong bump, which is the most extreme in our sample, suggests a high Fe abundance in the ejecta. A massive star that produces Fe must also show high abundance of O and Ca, as SN 1998bw did. However, SN 2022qml spectra do not show any O I features and only a small bump likely due to Ca NIR. Also, the shape of the light curve and its faster decline, compared to the other SNe in our sample, is suggestive of a different origin for this SN. It fact, it is possible that it was not a SN IIn but rather a Type Ia-CSM SN. This would be in line with what found by Leloudas et al. (2015), who showed that simulated thermonuclear SN spectra were consistently misclassified as SNe IIn once the underlying SN flux was or below with respect to the continuum. If this is the case, we probably missed the emergence of Si lines because the SN ejecta were still embedded in the CSM cocoon when the SN faded.
@@ 2.5 Host galaxies
The main characteristics of the host galaxies of the SNe in our sample are summarised in Table 2. The SN host galaxies as retrieved from PS1 in the band are shown in Fig. 9. Interestingly, all known hosts are small, probably dwarf galaxies.
@@
The case of SN 2022wed deserves some discussion, as the SN is close to three IR sources listed on NED (WISEA J072415.72+190456.3, WISEA J072415.05+190455.7, and WISEA J072415.95+190449.4) whose redshifts are unknown. Given the angular separation and assuming these galaxies have the same redshift of the SN, the radial distances from the centre of each galaxy would be 11, 15, and 17 kpc, respectively, which allows for the SN to belong to any of them. However, in the pre-explosion image produced by stacking PS1 observations between MJD , a faint source can be seen at the location of the SN (see Fig. 9, the bottom right panel). Its position is slightly offset from the SN, 0.01 arcsec in right ascension and 0.8 arcsec in declination, which, assuming it is at the same redshift of the SN, gives a radial distance of 1.7 kpc. Its FWHM is about 1.3 arcsec, comparable with a point-like source in the image we have. We apply a PSF-fit to the source and find an apparent magnitude mag, which, if we assume the same redshift of the SN, translates to an absolute magnitude of mag, similar to that of the host galaxies of the other SNe in the sample.
@@
Motivated by the presence of narrow [O III] and by the archival image of the host of SN 2021adxl, which shows a bright spot in correspondence with the SN position (see Fig. 9, the panel on the upper right), we attempted to use the line ratios of [N II] /H, [O III] /H, and [O II] /[O III] to derive an estimate of the oxygen abundance (which is assumed to trace the metallicity). We performed the measurements on the spectrum at +539, since it is the one where the SN contribution is smaller and the narrow lines due to the H II region are more evident. Unfortunately, we are only able to measure an upper limit for [N II] because the line cannot be resolved from the strong and broad H but the metallicity is consistent with a solar/subsolar composition according to the N2-versus-O3 calibrator. Based on a similar analysis but with spectra of better resolution, Brennan et al. (2023) report a subsolar (0.1 Z⊙) metallicity. Their relative flux measurements are consistent with ours but for H, which is three times lower than ours. This is not surprising considering that there is probably some residual contamination from the SN. For [N II] they have an actual measure 100 times smaller than our conservative limit. The metallicity measured by Brennan et al. (2023) is lower than the ones reported for SNe 2010jl and 2013L, both around 0.3 Z⊙ (Stoll et al., 2011; Taddia et al., 2020). This is consistent with regions of high star formation in low-mass galaxies such as the host of SN 2021adxl (Yates et al., 2012). For the other SNe, the hosts are too faint to properly distinguish the structure of the galaxy but, interestingly, the hosts of SNe 2021acya and 2022qml are also UV bright, which is indicative of a high specific star formation rate (SFR).
@@
In general, dwarf galaxies in the local Universe produce more stars per unit mass than massive galaxies and, furthermore, they seem to have a top-heavy initial mass function (IMF) that allows them to produce a higher fraction of massive stars than with standard Saltpeter IMF (Dabringhausen et al., 2012; Marks et al., 2012). This is consistent with the notion that strongly interacting SNe are massive stars from the higher end of the IMF and appear with higher frequency in dwarf galaxies.

### v2

@@ Abstract
Context. Massive stars experience strong mass-loss rates, producing a dense, H-rich circumstellar medium (CSM). After the explosion, the collision and continued interaction of the supernova (SN) ejecta with the CSM power the light curve through the conversion of kinetic energy into radiation. When the interaction is strong, the light curve shows a broad peak and high luminosity lasting for several months. Also the spectral evolution is slower, compared to non-interacting SNe. Energetic shocks between the ejecta and the CSM create the ideal conditions for particle acceleration and production of high-energy (HE) neutrinos above 1 TeV.
@@
The net result of an enormous (, Kiewe et al. 2012) mass-loss is a massive circumstellar medium (CSM) that is revealed by narrow lines in the spectrum after the explosion (e.g., Type IIn SNe, Schlegel 1990, 1996; Filippenko 1997) and contributes to the luminosity through interaction with the SN ejecta. The luminosity of interacting SNe is mainly powered by the conversion of kinetic energy into radiation in the shock of the ejecta with the CSM. It is expected that the shock is stronger when a shell ejection occurs shortly before the explosion (less than a few years). In fact, the closer the CSM shell is to the progenitor, the denser it is, and thus the SN ejecta colliding into it give rise to a more intense shock wave. When the CSM has a high density, the interaction can completely mask the internal power source to the extent that the underlying event can even be a non-terminal outburst (Vink, 2015), which could be mistaken for a faint core-collapse (CC) SN, or even a thermonuclear explosion (Silverman et al., 2013). The density of the CSM and its spatial distribution affect the shape of the light curve (Khatami & Kasen, 2023) and the strength of the emission lines, giving rise to asymmetries (Andrews et al., 2017) and peculiar line profiles such as double peaks (Andrews & Smith, 2018). Not only that, the mass and density of the CSM can also affect the shape of the light curve (Khatami & Kasen, 2023). Interestingly, Taddia et al. (2013) find that there does not appear to be a continuity of properties among SNe IIn but rather different subtypes, possibly hinting at the presence of different progenitors. Although the sample in Taddia et al. (2013) was quite small, this trend has been recently observed also with bigger samples of SNe IIn (Hiramatsu et al., 2024; Ransome & Villar, 2024).
Objects powered by strong interaction are rare. However, some extraordinary objects have been found in the past and present very well-sampled multiband light curves and spectra. A well-known SN that showed huge interaction in brightness and duration of the light curve is SN 2010jl (Smith et al., 2012b; Fransson et al., 2014; Ofek et al., 2014), a luminous Type IIn SN with a bright ( mag at peak) light curve that lasted more than 1000 and spectra dominated by H Balmer lines with symmetric, electron-scattering driven profiles. These characteristics are interpreted as signs of interaction with a massive (), H-rich CSM, possibly produced by an LBV progenitor (Fransson et al., 2014). LBVs and hypergiant stars have been proposed as progenitors of several SNe IIn (Gal-Yam et al., 2007; Gal-Yam & Leonard, 2009). Another example is SN 2013L (Andrews et al., 2017; Taddia et al., 2020), which also had an observed light curve lasting 1500 and peaking at around mag. The spectra are also slow-evolving and dominated by Balmer lines but the profile of the H line is significantly asymmetric, with a blue shoulder that is interpreted as due to the emergence of the shock wave not fully hidden by electron scattering (Taddia et al., 2020). We will use these two SNe as reference throughout the paper.
@@
Interaction is a powerful phenomenon that was proposed to power many SNe, among which, the superluminous SN class (SLSNe, Smith et al. 2007; Gal-Yam 2012), so as to explain why they are so luminous and with such slow evolution. The mechanism powering SLSNe is still debated, but the slow rise seems to favour a central engine (particularly, a magnetar, Kasen & Bildsten 2010) in a massive ejecta with long diffusion time. However, we caution that Hiramatsu et al. (2024) showed that there is no clear transition between SNe IIn and SLSNe and that the arbitrary cut should be removed.
In some cases, the CSM distribution may be strongly asymmetric. If it is also particularly close to the progenitor, the ejecta can quickly and completely engulf it. Narrow lines are then no longer visible in the spectra and the broad lines only show P-Cygni profiles typical of an expanding photosphere, but interaction continues within the inner ejecta, providing energy to the light curve. In these cases, it is difficult to determine whether the luminosity is fully due to interaction or to a central engine such as a magnetar, but the extreme energy and duration of these events seem to point toward a combination of the two phenomena (Kangas et al., 2022; Pessi et al., 2023; Salmaso et al., 2023).
@@ 2 The sample
Our goal is the selection of strongly-interacting SNe among all the SNe classified as IIn. An example of such objects found in the literature is SN 2010jl. These transients are the most appealing candidates for emitting HE neutrinos, although other less powerful objects cannot be ruled out. The SNe that correspond to our criteria in the literature are quite rare and often with little data, especially at very late phases. In this work, we build a sample of well-observed strongly-interacting SNe to analyse their properties. The criteria we use to identify the new transients may be useful in the future to enlarge this sample.
The sample was built selecting newly-discovered transients announced through services such as AstroNotes111https://www.wis-tns.org/astronotes and Astronomer’s Telegrams222https://astronomerstelegram.org that met our selection criteria. We evaluated both the classification spectrum and the light curve and selected SNe classified as Type IIn, with a relatively fast rising time among the luminous SNe ( days), and high but not extreme luminosity (). These criteria tend to exclude SLSNe, which are at least one order of magnitude more luminous ( or mag) and tend to have a longer rise time to peak ( d) (Lunnan et al., 2015). Despite the claims in Hiramatsu et al. (2024), the debate on the mechanism powering SLSNe is not fully settled, therefore, we decided to exclude them altogether from the sample, although this does not mean that they cannot be powered by interaction or produce HE neutrinos. Also, a preference was given to nearby events (typically within redshift , so that a more complete follow-up could be ensured). Following our procedure, we selected 11 targets in three years, which are listed in Tab. 3, together with their coordinates, redshift, distance modulus, and galactic absorption. We also report some notable parameters derived from the pseudo-bolometric light curve (for details on the computation, see Sect. 3), namely: the rise time , defined as the time between the last non-detection and the first peak; the peak luminosity ; the integrated luminosity in the first 200 days . These parameters will be used as criteria to ensure that the sample is not polluted by objects that show interaction at very late times or that underwent brief bursts of interaction over short periods of time. The case of SN 2021adxl is a bit peculiar since its poorly-constrained rise time is days, but given the light curve shape and luminosity, very similar to SN 2010jl (Fransson et al., 2014), it is likely that the explosion date was shortly before the peak. Therefore, we elected to keep it in the sample.
@@
As we will detail in the following, it turned out that not all SNe we selected with our criteria proved to evolve as expected and 7 out of 11 candidates were excluded from more detailed analysis for not showing signs of strong interaction. Their light curve was generally dimmer than the four strongly-interacting SNe that remained in the sample and faded faster. This can clearly be seen in Fig. 1, where we show their pseudo-bolometric light curve built using only optical filters (see Sect. 3 for the computation).
@@
To discriminate the transients that display strong interaction, we used the measured luminosity integrated over the first 200 and imposed that this parameter . This criterion is determined a-posteriori to identify the objects with prolonged, strong interaction. In fact, with longer integration times, the contamination arises from objects undergoing prolonged but mild interaction (e.g., SN 2005ip, Taddia et al., 2013, ). Conversely, with shorter integration times there is significant contamination from objects experiencing strong interaction for a limited period (e.g., SN 1998S). Our parameter is a compromise between the two cases and ensures that all the selected SNe are strongly-interacting. The excluded SNe had a lower energy display in the first 200 days compared to the selected ones. This can be seen from Fig. 2, where we also add the values calculated for the strongly-interacting SN 2010jl, the mildly-interacting SN 1998S, and the normal Type IIP SN 2004et (Sahu et al., 2006) as reference for different amounts of interaction. We also add the seven SNe IIn in the sample by Taddia et al. (2013): SNe 2005ip, 2005kj, 2006aa, 2006bo, 6006jd, 2006qq, and 2008fq. These prove to be polluters, since, although they show signs of interaction in the spectra, they are less luminous and their light curve has a shorter duration, thus implying the presence of less interaction. An intermediate case is that of SN 2021gci, which has significantly more energy than the others, and with a more luminous first peak, albeit still one order of magnitude below the selected transients. On the other hand, SNe 2022iaz and 2022owx have more luminous peaks but faded too fast, providing a small total amount of energy.
@@
This exercise denotes that a single spectrum and the early light curve is not enough to discern between strongly-interacting and normal interacting SNe, thus causing the sample to be contaminated with objects that have to be rejected later. The very minimum information required to identify a strongly-interacting SN includes at least photometric coverage of 2/3 months and a couple of good resolution spectra with enough signal-to-noise (S/N) at different epochs to check the evolution. Ideally, one should have a well-sampled light curve for the first 200 days to properly distinguish between strongly-interacting SNe and SNe with low interaction based on the energy input.
@@
The four SNe selected here that fulfill our criteria are then 2021acya, 2021adxl, 2022qml, and 2022wed. In the following, we will describe each object in more detail. We note that a paper presenting detailed observations for one of the targets in our list, SN 2021adxl, was recently published (Brennan et al., 2024). We will refer to their observations and results in the relevant sections.
@@ 2.1 SN 2021acya
SN 2021acya was discovered on October 2021 (MJD 59518.029) in the band by the Asteroid Terrestrial-impact Last Alert System (ATLAS, Tonry et al. 2018) at a magnitude of 18.124, while the last non-detection was only two days prior (Tonry et al., 2021). As the explosion epoch, then, we took MJD . It was then classified as SN IIn on November 2021 (Ragosta et al., 2021).
@@ 2.3 SN 2022qml
The spectral evolution of SN 2022qml is shown in Fig. 7. The spectrum at +24 has a blue continuum, implying a high temperature, and the only notable feature is the narrow H on top of a very broad but shallow emission. The spectrum at +49 starts to show a distinctive blue bump due to a forest of Fe II lines and a broad, asymmetric H with a boxy profile smoothed by electron scattering wings. We also identify the narrow line of [Fe X] and, possibly, [Ne V] , [Fe XI] and a broader emission that could be due to a blend of coronal [Fe IV] with [Ca V] . The absorption in the P-Cygni profile of the narrow H is also clearly visible, which allows for measuring the velocity of the progenitor wind 444We refer to it as wind for simplicity but the CSM could have been produced also by eruption or binary interaction. This is true throughout the paper. (). This feature is not identifiable in the other spectra of this object due to their lower resolution. At +56 the only prominent features are Balmer lines, [Fe X] , and He I . After that, the high ionisation lines disappear while the spectra remain dominated by the H, until phase +111, when the Ca II NIR emerges with a broad, symmetric emission. A zoom on the H region of this spectrum is also plotted in Fig. 6. The shape of the blue shoulder is similar to SNe 2021adxl and 2013L but it is broader, indicating a higher shock velocity.
The strong blue bump of SN 2022qml is reminiscent of a feature seen in Type Ia-CSM SNe (e.g., SN 2002ic, Kotak et al. 2004) but also in Type Ic SNe (e.g., SN 1998bw Galama et al. 1998; Kulkarni et al. 1998). As we mentioned, the presence of the blue bump is usually attributed to a forest of Fe II lines. In the case of SN 2022qml, the strong bump, which is the most extreme in our sample, suggests a high Fe abundance in the ejecta. A massive star that produces Fe must also show high abundance of O and Ca, as SN 1998bw did. However, SN 2022qml spectra do not show any O I features and only a small bump likely due to Ca NIR. Also, the shape of the light curve and its faster decline, compared to the other SNe in our sample, is suggestive of a different origin for this SN. It fact, it is possible that it was not a SN IIn but rather a Type Ia-CSM SN. This would be in line with what found by Leloudas et al. (2015), who showed that simulated thermonuclear SN spectra were consistently misclassified as SNe IIn once the underlying SN flux was a fraction or below with respect to the continuum. If this is the case, we probably missed the emergence of Si lines because the SN ejecta were still embedded in the CSM cocoon when the SN faded.
@@ 2.5 Host galaxies
The main characteristics of the host galaxies of the SNe in our sample are summarised in Table 2. The SN host galaxies as retrieved from the Panoramic Survey Telescope and Rapid Response System (Pan-STARRS, PS1, Chambers et al. 2019) in the band are shown in Fig. 9. Interestingly, all known hosts are small, probably dwarf galaxies.
@@
The case of SN 2022wed deserves some discussion, as the SN is close to three IR sources listed on NED (WISEA J072415.72+190456.3, WISEA J072415.05+190455.7, and WISEA J072415.95+190449.4) whose redshifts are unknown. Given the angular separation and assuming these galaxies have the same redshift of the SN, the radial distances from the centre of each galaxy would be 11, 15, and 17 kpc, respectively, which allows for the SN to belong to any of them. However, in the pre-explosion image produced by stacking PS1 observations between MJD , a faint source can be seen at the location of the SN (see Fig. 9, the bottom right panel). Its position is slightly offset from the SN, 0.01 arcsec in right ascension and 0.8 arcsec in declination, which, assuming it is at the same redshift of the SN, gives a radial distance of 1.7 kpc. Its Full-Width-at-Half-Maximum (FWHM) is about 1.3 arcsec, comparable with a point-like source in the image we have. We apply a Point-Spread-Function (PSF) fit to the source and find an apparent magnitude mag, which, if we assume the same redshift of the SN, translates to an absolute magnitude of mag, similar to that of the host galaxies of the other SNe in the sample.
@@
Motivated by the presence of narrow [O III] and by the archival image of the host of SN 2021adxl, which shows a bright spot in correspondence with the SN position (see Fig. 9, the panel on the upper right), we attempted to use the line ratios of [N II] /H, [O III] /H, and [O II] /[O III] to derive an estimate of the oxygen abundance (which is assumed to trace the metallicity). We performed the measurements on the spectrum at +539, since it is the one where the SN contribution is smaller and the narrow lines due to the H II region are more evident. Unfortunately, we are only able to measure an upper limit for [N II] because the line cannot be resolved from the strong and broad H but the metallicity is consistent with a solar/subsolar composition according to the N2-versus-O3 calibrator. Based on a similar analysis but with spectra of better resolution, Brennan et al. (2024) report a subsolar (0.1 Z⊙) metallicity. Their relative flux measurements are consistent with ours but for H, which is three times lower than ours. This is not surprising considering that there is probably some residual contamination from the SN. For [N II] they have an actual measure 100 times smaller than our conservative limit. The metallicity measured by Brennan et al. (2024) is lower than the ones reported for SNe 2010jl and 2013L, both around 0.3 Z⊙ (Stoll et al., 2011; Taddia et al., 2020). This is consistent with regions of high star formation in low-mass galaxies such as the host of SN 2021adxl (Yates et al., 2012). For the other SNe, the hosts are too faint to properly distinguish the structure of the galaxy but, interestingly, the hosts of SNe 2021acya and 2022qml are also UV bright (as per the magnitudes reported in NED), which is indicative of a high specific star formation rate (SFR).
@@
In general, dwarf galaxies in the local Universe produce more stars per unit mass than massive galaxies and, furthermore, they seem to have a top-heavy initial mass function (IMF) that allows them to produce a higher fraction of massive stars than with standard Saltpeter IMF (Dabringhausen et al., 2012; Marks et al., 2012). Interestingly, SNe IIn in general do not tend to follow the star formation as traced by the H emission (Habergham et al., 2014), possibly indicating multiple progenitor scenarios (Ransome et al., 2022). However, luminous SNe IIn do occur more frequently in metal-poor environments with young stellar populations (Moriya et al., 2023). This is consistent with the notion that strongly interacting SNe are massive stars from the higher end of the IMF and appear with higher frequency in dwarf galaxies.